### v2

# Visual collective behaviors on spherical robots

Abstract
The implementation of collective motion, traditionally, disregard the limited sensing capabilities of an individual, to instead assuming an omniscient perception of the environment. This study implements a visual flocking model in a “robot-in-the-loop” approach to reproduce these behaviors with a flock composed of 10 independent spherical robots. The model achieves robotic collective motion by only using panoramic visual information of each robot, such as retinal position, optical size and optic flow of the neighboring robots. We introduce a virtual anchor to confine the collective robotic movements so to avoid wall interactions. For the first time, a simple visual robot-in-the-loop approach succeed in reproducing several collective motion phases, in particular, swarming, and milling. Another milestone achieved with by this model is bridging the gap between simulation and physical experiments by demonstrating nearly identical behaviors in both environments with the same visual model. To conclude, we show that our minimal visual collective motion model is sufficient to recreate most collective behaviors on a robot-in-the-loop system that is scalable, behaves as numerical simulations predict and is easily comparable to traditional models.
[…]
  • October 6, 2024
Keywords: Flocking, robot-in-the-loop, Collective motion, Milling, Swarming, Optic Flow, Visual information.
1 Introduction
Biological collective motion has inspired efforts to understand, replicate, and ultimately utilize on artificial agents. Nearly 40 years ago, Reynolds et al.,[1] proposed that a set of simple, individual-level rules could explain collective motion. These rules include (i) attraction, the desire to remain close to other individuals; (ii) alignment, the tendency to move in the same direction as the group; and (iii) avoidance, the willingness to avoid collisions with others. Vicsek et al.,[2] later introduced a model based on statistical analysis, which averaged the velocities of individuals within a defined proximity of the individual and accounted for individual tendencies as stochastic noise. Couzin et al.,[3] expanded upon these ideas by combining the rules defined by Reynolds with the area of interaction defined by Vicsek. Authors propose that each rule have an area of interaction: the closest avoidance, the following alignment and the biggest one attraction. These 3 rules effectively reproduce in simulation three collective behaviors observed in nature, swarming (no orientational order), schooling (highly polarized movement), and milling (circular, swirling movement). Models that follow these guides are now known as “3A models”[4, 5, 6].
Most of available approaches to 3A-models rely on artificial tools to describe or enforce the rules, such as inter-individual distance[7, 8, 9], global velocity[10, 11], or absolute communication among individuals[12, 13]. However, these methods do not accurately reflect the way animals interact with their environment or others. To address this gap, researchers have explored biologically plausible (“bio-plausible”) ways of describe the environment[14, 15, 16, 17], with vision proving to be the most effective and widespread one[18, 19, 20, 21, 22]. Several vision-based models have been tested on robotic platforms[23, 24], but both of these either use vision merely as a selective mechanism[25, 26] or lack alignment with the fundamental principles of the 3A models[27].
Castro et al. (2024) [28] introduced a model that utilizes only visual information while maintaining the traditional 3A model structure. This model effectively replicates the phases of collective motion described by Couzin et al., by employing a visual mechanism known as optic flow. Optic flow, which captures the apparent angular motion of the environment due to the relative movement between an observer and its surroundings, has been validated as a crucial navigational tool in nature. Optic flow magnitude is used by insects during most flight maneuvers [29, 30] as well as in fish and birds [31, 32]. Furthermore, several robotic platforms have already incorporated small sensors dedicated to measuring optic flow, highlighting its potential as a key mechanism for both biological understanding and innovation[33, 34, 35, 36, 37].
Building on these findings, this study proposes a “robot-in-the-loop” approach to collective motion that utilizes only visual information, implemented on a educational robotic platform. Unlike previous models, this approach does not rely on global information or complex communication networks. Instead, it introduces a visual anchor to account for environmental boundaries, thereby eliminating the need to model wall interactions explicitly. Experimental results and numerical simulations demonstrate that this method successfully reproduces several collective motion behaviors on a flock of 10 independently controlled robots.
The article is organized as follows: Section 2.1 provides an overview of the challenges of a robotic implementation of 3A-models, with the example of the visual-based collective motion model proposed by Castro et al. (2024) [28] (Fig.1). Section 2.2 describes the robotic platform and the model modifications for implementing on the 10 independently controlled robots including the introduction of the visual anchor concept. Section 2.4 detail the implementation of a feature matching optic flow and visual collective motion rules. Section 2.5 describes the supplementation on original model and summarizes robotic setup, while Section 3 presents the results from simulations and robot-in-the-loop experiments. Finally, Section 4 concludes with a discussion of the results and directions for future research.
2 Methods
Figure 1: Model collective behaviors from optic flow and retinal cues [28]. a) View of the flock positions.b) Representation of the flock positions after recognition at time step $t$. Each $i$-particle is described by its XY-coordinates and unit heading vector $\mathbf{e}_{i}$, with the heading angle $\theta_{i}$. c) Point-of-view of particle 0 showing the retinal relative positions of each particle. The arrows represent the perceived optic flow, and the white arrow indicates the apparent velocity ($\mathbf{V}_{i,k}$). Here, we only see the azimuthal component of the apparent velocity $\mathbf{V}_{i,k}$ by the particle i=“0” of retinal object “k”. d) $\mathcal{V}(\phi)$ binary panorama representation of the field of view of particle 0. e) Apparent distance $\mathcal{R}_{i}(\phi)$ of particle “0”. f) Optic flow $\mathcal{O}_{i}(\phi)$ perceived by particle 1. g) Optic flow divergence $\mathcal{D}_{i}(\phi)$ perceived by particle 1.
2.1 Robotic Implementation challenges
Flocking models that are based on individual rules of attraction, avoidance and alignment, or better known as 3A-models, present several challenges when being implemented on a robotic platform. The biggest two are related to how individuals are modeled. The assumption of most models is that each individual is a self propelled particle, traditionally moving on a plane. Yet most robots are particles or restrict their movement only to a plane. This brings forth the first challenge, which robot kinematics should be chosen? It should balance following the assumptions of the model, sensing capabilities, computational capacity, size, among others. By exiting the numerical environment, the experimental environment are not infinite anymore, so the collective behavior should be confined which is a second challenge. Without this considerations flocking on robots would be as direct as a raw implementation of a established model, such as Castro et al., (2024) [28] on a the selected robotic platform (as seen on Fig. 1). The following sections will present our solution to both challenges.
2.2 Robotic setup
This section will section will present our robot setup to a robot-in-the-loop visual flocking model. We propose to use an underpowered, commercially available robotic platform that offloads the sensing capabilities to a base station but independently uses the information from its binary visual map to execute the model presented in the previous section.
Figure 2: Robotic platform. a) Sphero Bolt®wireless educational robots. b) Example of the robots’ movement, c) Heading control by differential rotation on the two opposite inner wheels, d) Induced nose-up pitch during forward displacement.
This robot-in-the-loop solution utilizes an existing educational robotic platform, the Sphero Bolt®(Fig. 2a). These are small, spherical robots that can be controlled in real-time via Bluetooth®and an educational app, where users can set a heading value, desired velocity, LED states, and read the onboard sensors’ values. Additionally, several open-source projects integrate these robots with different programming environments such as ROS, Python, or MATLAB®.
Each robot is equipped with several sensors, including an Inertial Measurement Unit (IMU) comprising a 3-axis accelerometer, gyroscope, and magnetometer, and a light sensor. It also features four infrared (IR) proximity sensors on its sides, two addressable RGB LEDs (one in the front and one in the back), and an 8x8 RGB LED matrix on top.
The Sphero’s movement is achieved via a self-balancing two-wheel enclosed system, with two non-actuated wheels maintaining constant contact with the enclosing sphere. This dual-motor design allows for stationary turning; however, while moving, the turning is achieved by differential motion (Fig. 2b-c). The locomotion generates a pitch displacement depending on the speed and the friction coefficient between the sphere’s encapsulation and the surface (Fig. 2d). This pitch is partially managed by the robot’s dual-MCU configuration: one MCU handles the communication protocol and high-level instructions, while the other manages motor control to ensure the robot achieves the desired heading and speed, along with low-level sensor readings.
Figure 3: Experimental setup. a) Picture of the experimental setup, b) Example of the images taken by the camera (Daheng Imaging MER2-230-168U3C camera with 4mm c-mount lens), c) Whole robot body image segmentation, d) Individual light sources and heading recognition, e) Individual robot identification and virtualization.
To address the lack of an onboard camera and the inability to embed one into the robots, we set up a bird’s-eye camera above the experimental area. This is an industrial RGB global shutter camera, the Daheng Imaging MER2-230-168U3C, with a frame rate of 168 FPS, a resolution of 1920x1200, and support for interchangeable c-mount lenses (Table 1). The selected lens is a 4mm, 1/8”, F2.0-12 non-distorting lens, which allows the camera to be mounted 2.3 m above the robots, covering a 2.7x2.7 m working area (approximately 38 times the length of the robots) (Fig. 3a).
Shutter time 25 ms
Pixel Bit Depth 10 bit
Digital gain 5 dB
Pixel Data Format Bayer RG10
Lens f-stop 11
Interface Python 3.10
Table 1: Camera Settings.
Each robot is assigned an individual color pattern that enables identification of the robot’s front and back. To maintain light connectivity between the front and back, the back half of the LED matrix is set to 50% bright white light. This configuration provides 16 possible combinations for identifying each robot, with the potential for expansion using additional colors. The bird’s-eye camera allows the isolation of each light source on top of the robots, preventing light bleeding between robots when they are in close proximity (Fig. 3b). This is achieved by capturing a 1920x1200 RGB image and performing whole-robot segmentation. This segmentation treats the three light sources as a single entity, allowing calculation of the center’s XY-coordinates, which approximate the robot’s center. Knowing the radius of the robots in pixels allows for cropping of $N$ images from these coordinates, each isolating a single robot. The cropped images are further segmented to identify the front and back LEDs, as well as the LED matrix. The larger area generated by the two light sources enables front and back recognition (Fig. 3d). Additionally, by correlating the color from this segmentation, each robot can be uniquely identified. The final step is to compute the robot’s heading by tracing the vector between the center positions of the front and back LEDs. This process results in the identification, XY-coordinates, and heading of each robot (Fig. 3e).
The next step, after identifying each robot and determining their location and heading, is to reconstruct the binary panorama of each robot.
2.3 Anchoring the collective behavior
Figure 4: Binary panorama reconstruction. a) 2D virtual representation of the binary panorama, the red goal represents the desired location of the flock. b) Mixed reality point of view of the 0-particle with the omniscient red goal. c) Graph representation of all $\mathcal{V}(\phi,t)$. $\mathcal{V}(\phi,t)$ is a binary panorama representation of the field of view of particle i where the anchor is represented a different channel (red). This graph can be stored as a tuple list (rising and falling edges).
In the binary panorama reconstruction, we introduce a new component to the original model [28]: a visual anchor. The inclusion of an anchor bounds the flock to a place and prevent infinite expansion. The anchor is an augmented reality (AR) solution that enables each robot to see the desired center point (Fig. 4a). This anchor is omnisciently visible, regardless of any obstacles between the line of sight and the anchor location (Fig. 1b). It is a new visual channel that uses the same binary representation as the visual surroundings (Fig. 1c). The binary panorama is represented in Python as an ordered tuple list of every retinal object’s rising and falling edge, with the last position always being the goal tuple. This process is performed independently for each particle without sharing information between particles.
2.4 Robotic Visual Flocking Model
Avoidance effect ($\mathcal{A}$) is here introduced in the attraction term ($\omega_{\odot}$). Our previous study [28] did not include any avoidance because most collective motion models are based on the assumption that all particles have some extra room on the z-axis, allowing them to avoid collisions with additional depth. This assumption is primarily based on how fish flocking data is gathered, where the fish can swim above or below each other in a few centimeters of water [38]. However, for any robotic implementation, obstacle or collision avoidance is of critical importance. Fortunately, avoidance and attraction have similar characteristics, being proportional instantaneous responses to the optical size and location of a retinal object. The key difference is in the direction of the response: while attraction is positive, avoidance is negative. Therefore, avoidance is implemented by providing a new piecewise constant function ($\mathcal{A}$), where attraction is inverted in situations where a particle should avoid a retinal object. This new function depends on the width of the retinal object, specifically the difference between its rising and falling edges ($\Delta\phi$). This being a indirect measurement of the apparent distance ($\mathcal{R}(\phi,t)$). We choose $\mathcal{A}(\Delta\phi)=-5$ for $\Delta\phi<\pi/10$, which is equivalent to the retinal object being closer than three particle radii ($\mathcal{R}<3a$). In all other cases, $\mathcal{A}(\Delta\phi)=1$. The blind spot remains unchanged, giving more weight to the retinal object in front of the particle, regardless of this new function.
Alignment ($\omega_{\parallel}$) remains unchanged with respect to previous study [28], considering that the divergence/convergence of the optic flow already accounts for retinal objects approaching the focal object.
The optic flow for each retinal object can be decomposed in two components: the azimuthal in its center ($\overline{\mathcal{O}}$) and its radial divergence ($\mathcal{D}$). The former represents the apparent displacement of retinal features, while the latter represents changes in the size of those features. Divergence is a decrease in size, and convergence is an increase. With a 1D representation, these features are the rising and falling edges of the retinal object, akin to a feature-matching optic flow calculation.
Calculating optic flow as a feature-matching operation heavily relies on accurately identifying the same features across consecutive time steps. The model accomplishes this by finding the gradient of the Euclidean distance of the retinal object position and size between consecutive time steps. Regardless of flock size, there are three possible cases. If, in both time steps, there are the same number of retinal objects, the minimum of the gradient will be continuous and located on the main diagonal of a square matrix. When the number of retinal objects differs, it indicates either a separation or merge of retinal objects. In this case, there is no match, and the retinal objects are ignored in the calculation.
The apparent velocity vector $\mathbf{V}_{ik}$ in polar coordinates is $\mathbf{V}_{ik}=[-\mathcal{D}_{i},\overline{\mathcal{O}}_{i}]\cdot\mathcal{R}_{i}/U$ with the heading vector $\mathbf{e}_{i}=[\cos\phi,-\sin\phi]$. The supplemented visual model equations for the collective behavior are presented in Eqs. 1a-$f$. Note that the noise $k_{\eta}$ and the time delay $\tau$ are set to 0 for robot-in-the-loop experiments, as they are parameter modeling real-world noise and uncertainty.
$\dot{\mathbf{x}}_{i}(t)=U\mathbf{e}_{i}(t),$ (1a)
$\dot{\theta}_{i}(t)=k_{\odot}\,\omega_{\odot}(t)+k_{\parallel}\,\omega_{\parallel}(t)+k_{c}\,\omega_{c}(t)+k_{\eta}\,\eta(t),$ (1b)
[…]
$\frac{\mathbf{e}_{i}(t)\times\mathbf{V}_{ik}(t)}{\mathcal{R}_{i}(t)}=-\mathcal{D}_{i}(\phi,t)\sin(\phi,t)+\overline{\mathcal{O}}_{i}(\phi,t)\cos(\phi,t).$ (1f)
Lastly, the robotic platform introduces several critical processing factors – Bluetooth, image processing, internal Spheros control command, … – that affect the overall dynamical response of the robots and its collective behavior as a whole (see Supplementary Information for more details). These extra factors results in a time delay $\tau$ with a median of 600ms. The overall functional diagram of the collective motion model is presented in Fig. 5. This model is executed independently and simultaneously for all robots; it does not involve communication between robots, sharing information, and only uses visual data. The same block could be utilized on a robotic platform using LiDAR, onboard cameras, touch sensors, ultrasound, or any other sensor that provides a binary visual panorama as output. All the information used by this model is self-referred and self-contained; there is no global reference frame, coordination, or knowledge of the flock or the environment.
Figure 5: Functional diagram of the collective motion algorithm with robots-in-the-loop.
2.5 Robots-in-the-loop computer architecture
We propose a robots-in-the-loop flocking that consists of a set of $N$ Sphero® Bolt robots, which are devoid of visual sensor and are controlled via Bluetooth®. Their positions and headings are calculated by processing the images from an industrial 1920x1200 RGB, 164 FPS overhead camera. From all their positions, a binary visual map is recreated. This visual map serves as the input for the collective motion model. This process is shown in Fig. 6.
Figure 6: General Block diagram. The N flocking robots are seen by the 1920x1200px, 164fps camera. This image is acquired and processed to find the XY-coordinates, heading and colors for all $N$ robots. With these coordinates each robot can be identified. The XY-coordinates of the Anchor can be set at any point of the image to be integrated to the identification of the robots. All these coordinates are processed to reconstruct the binary panorama for each of the spherical robots. Each binary Panorama is the input of independent Collective motion model that process the visual panoramic information. All these models are ran in parallel and their output in term of heading are sent to each robot via a Bluetooth interface.
There are two main Python programs to handle the implementation of these blocks. The first one (administrator.py) manages the camera communication, data extraction, image processing, object identification, and binary panorama reconstruction (Fig. 3). The output is transferred to $N$ independent routines named PyRobot.py (the second program), which take a single binary panorama, perform the collective motion model, manage the Bluetooth connection, and send the desired heading and velocity (Fig. 5). Each robot receives the inputs (heading and velocity) and moves accordingly. Each PyRobot.py is a routine that can be executed either locally or remotely from the Administrator. This setup allows for scaling the flock size to the maximum number that can fit within the camera’s recognition area. Note that this scaling is hardware-constrained by the Bluetooth interface capabilities of the computer. For Windows®machines, the theoretical maximum is seven devices per machine; in Linux, it is seven per interface. However, it is recommended to have a maximum of five on the same machine managing the camera and six per interface on independent machines.
2.6 Metrics
To quantitatively distinguish between different dynamical phases, we introduce three global order metrics: polarization ($P$), milling ($M$), and opacity ($O$) [39, 40]. These metrics are defined as follows:
$P(t)$ $=$ $\left\|\overline{\mathbf{e}_{i}(t)}\right\|,$ (1b)
$M(t)$ $=$ $\left\|\overline{\mathbf{y}_{i}(t)\times\mathbf{e}_{i}(t)}\right\|,$ (1c)
[…]
where the overbar denotes an average over all individuals, and the unit vector $\mathbf{y}_{i}(t)=\frac{\mathbf{x}_{i}(t)-\overline{\mathbf{x}_{i}}(t)}{\|\mathbf{x}_{i}(t)-\overline{\mathbf{x}_{i}(t)}\|}$ points towards particle $i$ from the center of mass. All three metrics range within the interval $[0,1]$. The polarization $P(t)$ measures alignment: $P(t)=0$ corresponds to particles pointing in all directions, while $P(t)=1$ represents a perfectly aligned school. The milling $M(t)$ quantifies the normalized angular momentum: $M(t)=0$ corresponds to a straight-line formation, and $M(t)=1$ represents perfect milling. The opacity $O(t)$ measures the “occupancy” of the visual fields: $O(t)=0$ indicates no object in the visual field, and $O(t)=1$ signifies that the entire visual field is obscured.
3 Results
3.1 Simulation of supplemented visual model
To make the problem dimensionless, we chose $a=1$ and $U=1$. We also kept the blind angle parameter and noise strength with values that lead to the most phase reproduction: $\epsilon=1$, $k_{\eta}=0.01$. With this approach, three dimensionless parameters remain: the strengths of anchor attraction and alignment, $k_{c}$, $k_{\odot}$, and $k_{\parallel}$. The anchor is a parameter aiming for a contained robotic experimentation, set practically to the minimum value that ensures the robots do not reach the arena’s border when no other law is in effect, $k_{c}=0.005$.
The visual original model has been supplemented with three major changes impacting the phase distribution and the metric on the parameter space $(k_{\odot},k_{\parallel})$:
  • the inclusion of the avoidance rule ($\mathcal{A}(\phi)$) (Fig. S7),
  • the addition of an anchor law ($\omega_{c}$) (Fig. S8),
  • the introduction of a time delay ($\tau$) (only in simulation) to reflect the time delay between the desired heading and a perceived change in heading while operating the Spheros robot, (Fig. 7).
Figure 7: Illustration of the phases observed in simulation for the supplemented visual model with the addition of avoidance, anchor, and time delay as defined in Eqs. 1a, with $\tau=2.4[a/U]$. Three phases are observed when varying the other two parameters: a) Swarming ($k_{\odot}=0.25$, $k_{\parallel}=0.04$) (see SVideo 1); b) Bistable ($k_{\odot}=0.2$, $k_{\parallel}=0.25$) (see SVideo 2); c) Milling ($k_{\odot}=0.07$, $k_{\parallel}=0.16$) (see SVideo 3).
We now compare the original model with each consecutive addition by exploring the parameter space $(k_{\odot},k_{\parallel})\in[0,0.3]\times[0,0.3]$ on a flock of 10 individuals. For each parameter set, we ran 10 simulations over a long duration ($\Delta t=5000$). The mean values of $P$, $M$, and $O$ were determined by averaging over the last 1000 time units to ensure that transients do not influence the results. The outcomes of these simulations are synthesized in phase diagrams (Fig. 8).
We maintain the identification of the collective phases from the values of the polarization $P$ and milling $M$ parameters: schooling when $P>0.5$ and $M<0.5$, milling when $P<0.5$ and $M>0.5$, and swarming when $P<0.5$ and $M<0.5$.
Figure 8: Simulation Phase Diagrams for 10 individuals. The dotted line is $P=0.5$, the dashed line is $M=0.5$, and the solid line is $O=0.3$. A) Color code, b) Original visual model [28], c) Supplemented visual model with avoidance term, $\mathcal{A}(\mathcal{R}(\phi,t))$, d) Supplemented model with avoidance and anchor term, $\omega_{c}(t)$, e) Supplemented visual model with avoidance, anchoring and perceived time delay ($\tau$=2.4[a/U]).
3.2 Phase diagrams of the supplemented model in simulation
We now compare the model with each consecutive addition and the robotic implementation, exploring the parameter space $(k_{\odot},k_{\parallel})\in[0,0.3]\times[0,0.3]$ for a flock of 10 independently controlled robots. For the robotic implementation, the dimensionless parameters are set to match the radius in pixels and velocity in radius per second, resulting in $a=1$, with the other parameters remaining the same ($\epsilon=1$, $k_{\eta}=0$, and $k_{c}=0.005$). For each parameter set, we ran 5 robotic experiments over a long duration (240sec or $\Delta t\approx 8500$). The mean values of $P$, $M$, and $O$ were determined by averaging over the last 1000 time units to ensure that transients do not influence the results. The outcomes of these experiments are synthesized in phase diagrams (Fig. 9a). An example of swarming is shown in Fig. 9b, of partial schooling for a bistable phase in Fig. 9c, and of milling in Fig. 9d.
The phase diagram of the original model [28] showcases the reproduction of the main three phases (swarming, schooling, and milling), with the area above $k_{\odot}>0.15$ and $k_{\parallel}>0.15$ being bistable (Fig. 8a). A bistable phase oscillates between milling and schooling (Fig. 8b). By adding repulsion to the model, this bistable phase is removed, but the area where milling appears shrinks to roughly half of the original while the final opacity of the flocks is reduced. Opacity is an inverse metric of flock density, and this reduction aligns with the expected effect of adding avoidance.
[…]
Fig. 8c shows the introduction of the anchor, which eliminates the schooling phase while the swarming phase area remains unchanged. This change is attributed to the turning effect generated by the anchor. When all particles are in a line, the front particles turn faster than those at the back, resulting in a turn that eventually stabilizes into a circle and a very stable milling pattern.
Adding a time delay ($\tau=2.4[a/U]$) in the supplemented visual model (Fig. 8d) causes the swarming phase to expand to values above $k_{\odot}=0.17$ or when $k_{\parallel}<0.08$. For low attraction ($k_{\odot}<0.03$), there is bistability, and the milling phase is reduced to an area for alignment above $k_{\parallel}=0.1$ and medium attraction ($0.170.06<k_{\odot}<$ 0.17).
3.3 Comparison with the phase diagram of the Spheros’ robots-in-the-loop
For the robotic implementation (see Fig. 9), we found that the milling phase remains fairly stable, appearing for high alignment ($k_{\parallel}>0.15$) across most attraction values ($k_{\odot}<0.16$). The swarming phase is constant for low alignment ($k_{\parallel}<0.06$) or high attraction ($k_{\odot}>0.17$).
The avoidance law, subject to the same time delay, results in some collisions that resolve over time, leading to higher opacity values. Similarly, for $k_{\odot}<0.01$, avoidance is too small, and the flock groups around the anchor without moving. This behavior is somewhat mitigated when there is enough alignment ($k_{\parallel}>0.15$), but the flock remains more compact than in the simulations.
4 Discussion
The collective motion model with the robotic platform reproduces several collective patterns such as swarming, milling, and a bistable phase. For the first time to the best of our knowledge, a visual flocking on spherical robots shows a phase diagram that matches almost perfectly that of the simulation with the same visual model (Fig. 8e vs Fig. 9a). The mean squared error (MSE) between the phase diagram images of robots-in-the-loop and supplemented visual model results is 2.48% (for more details see Fig. S2). Such implementation therefore bridges the gap between collective model simulation and real-world experiments, allowing for expansion to $N$ individuals with low complexity. A model based on visual information enables a mixed-reality approach where there can be interaction between the flock and the environment without adding complexity to the model or sensors available to each robot.
Future works with these spherical robots could include an explicit channel for avoidance also using optic flow, a moving anchor, a replacement of the anchor by another particle without adding a new channel, and an expansion of the flock size to 20 or 40 individuals for a closer-to-nature implementation of these visual collective models.
Acknowledgements
The participation of D.C. in this research was made possible by joint PhD grant from Aix-Marseille University.
D.C. and F.R. were also supported by Aix Marseille University and the CNRS (Life Science, Information Science, and Engineering and Science & technology Institutes).
The facilities for the experimental tests has been mainly provided by ROBOTEX 2.0 (Grants ROBOTEX ANR-10-EQPX-44-01 and TIRREX ANR-21-ESRE-0015).
Data Availability
Data acquired during the experiments and simulations are available online [41, 42]
Video Accessibility
Links to supplemental videos are provided in the supplementary information and can also be found at: youtube or [42].
Author contributions statement
D.C. performed the simulations, the robotic experiments, and data consolidation. C.E. & F.R. outlined the robot in the loop experiment and supervised its implementation. All authors: Proofread the manuscript.
See pages 1- of SuppData2.pdf

### v3

@@
 Another milestone achieved with by this model is bridging the gap between simulation and physical experiments by demonstrating nearly identical behaviors in both environments with the same visual model.
-To conclude, we show that our minimal visual collective motion model is sufficient to recreate most collective behaviors on a robot-in-the-loop system that is scalable, behaves as numerical simulations predict and is easily comparable to traditional models.
+To conclude, we show that our minimal visual collective motion model is sufficient to recreate most collective behaviors on a robot-in-the-loop system that be implemented using several individuals, behaves as numerical simulations predict and is easily comparable to traditional models.
 […]
-  • October 6, 2024
+  • 15 January 2025
 Keywords: Flocking, robot-in-the-loop, Collective motion, Milling, Swarming, Optic Flow, Visual information.
 1 Introduction
-Biological collective motion has inspired efforts to understand, replicate, and ultimately utilize on artificial agents.
-Nearly 40 years ago, Reynolds et al.,[1] proposed that a set of simple, individual-level rules could explain collective motion.
-These rules include (i) attraction, the desire to remain close to other individuals; (ii) alignment, the tendency to move in the same direction as the group; and (iii) avoidance, the willingness to avoid collisions with others.
-Vicsek et al.,[2] later introduced a model based on statistical analysis, which averaged the velocities of individuals within a defined proximity of the individual and accounted for individual tendencies as stochastic noise.
-Couzin et al.,[3] expanded upon these ideas by combining the rules defined by Reynolds with the area of interaction defined by Vicsek.
-Authors propose that each rule have an area of interaction: the closest avoidance, the following alignment and the biggest one attraction.
-These 3 rules effectively reproduce in simulation three collective behaviors observed in nature, swarming (no orientational order), schooling (highly polarized movement), and milling (circular, swirling movement).
-Models that follow these guides are now known as “3A models”[4, 5, 6].
-Most of available approaches to 3A-models rely on artificial tools to describe or enforce the rules, such as inter-individual distance[7, 8, 9], global velocity[10, 11], or absolute communication among individuals[12, 13].
-However, these methods do not accurately reflect the way animals interact with their environment or others.
-To address this gap, researchers have explored biologically plausible (“bio-plausible”) ways of describe the environment[14, 15, 16, 17], with vision proving to be the most effective and widespread one[18, 19, 20, 21, 22].
-Several vision-based models have been tested on robotic platforms[23, 24], but both of these either use vision merely as a selective mechanism[25, 26] or lack alignment with the fundamental principles of the 3A models[27].
-Castro et al.
-(2024) [28] introduced a model that utilizes only visual information while maintaining the traditional 3A model structure.
-This model effectively replicates the phases of collective motion described by Couzin et al., by employing a visual mechanism known as optic flow.
-Optic flow, which captures the apparent angular motion of the environment due to the relative movement between an observer and its surroundings, has been validated as a crucial navigational tool in nature.
-Optic flow magnitude is used by insects during most flight maneuvers [29, 30] as well as in fish and birds [31, 32].
-Furthermore, several robotic platforms have already incorporated small sensors dedicated to measuring optic flow, highlighting its potential as a key mechanism for both biological understanding and innovation[33, 34, 35, 36, 37].
-Building on these findings, this study proposes a “robot-in-the-loop” approach to collective motion that utilizes only visual information, implemented on a educational robotic platform.
-Unlike previous models, this approach does not rely on global information or complex communication networks.
-Instead, it introduces a visual anchor to account for environmental boundaries, thereby eliminating the need to model wall interactions explicitly.
-Experimental results and numerical simulations demonstrate that this method successfully reproduces several collective motion behaviors on a flock of 10 independently controlled robots.
+1.1 Related Work
+Biological collective motion has inspired extensive efforts to understand, replicate, and utilize it in artificial systems.
+Nearly four decades ago, Reynolds et al.
+[1] proposed that collective motion could emerge from simple, individual-level rules: (i) attraction, the tendency to stay close to others; (ii) alignment, the inclination to move in the same direction as neighbors; and (iii) avoidance, the need to prevent collisions.
+Vicsek et al.
+[2] introduced a statistical model where individuals align their velocities with neighbors within a given proximity, incorporating stochastic noise to reflect individual variation.
+Couzin et al.
+[3] expanded these ideas by integrating spatial interaction zones, with avoidance in the innermost region, alignment in the intermediate zone, and attraction in the outermost zone.
+These Attraction, Avoidance, Alignment rules (“3A”) effectively reproduce key collective behavioral phases observed in nature.
+We define these phases as: swarming [^1] Note, this ”swarming” phase is a different concept to swarm robotics.
+Here, the swarming phase is an emerging behavior that occurs when the group moves in a non-organized motion.
+(non-polarized cohesive motion), schooling (highly polarized motion, also known as flocking), and milling (circular motion).
+Models based on these principles are commonly referred to as “3A models”  [4, 5, 6, 7].
+Many “3A models” rely on artificial methods to enforce rules, such as inter-individual distances [8, 9, 10, 11], global velocity measurements [12, 13], or absolute communication networks [14, 15].
+However, such approaches often diverge from the mechanisms used by animals, prompting researchers to explore more biologically plausible (”bio-plausible”) models [16, 17, 18, 19].
+Vision-based mechanisms have proven particularly effective, as vision is the primary sensory modality for many species [20, 21, 22, 23].
+Vision-based collective motion models have previously been tested on robotic platforms [24, 25].
+Moreover, some authors have previously used only vision to successfully model collective motion [26, 27, 28, 29].
+Some implementations treated vision as a selective mechanism [30, 31] or do not fully adhere to the fundamental principles of the classical 3A model by not using its structure of “alignment”, “attraction” or “avoidance” rules [27, 32].
+Recently, Castro et al.
+(2024) [33] proposed a model that relies solely on visual information while maintaining the structure of 3A models.
+Their model replicates collective behaviors using “early vision” cues : (i) angular position, (ii) angular size in the field of view as well as (iii) optic flow, the visual perception of the apparent angular motion of the environment caused by relative movement.
+Optic flow is a key navigational tool in nature, used by insects during flight [34, 35] and by fish and birds [36, 37].
+Robotic implementations of optic flow have also demonstrated significant potential [38, 39, 40, 41, 42].
+1.2 Contribution
+This study proposes a robot-in-the-loop approach to collective motion that relies exclusively on these “early vision” cues using educational spherical robots.
+Unlike previous models, this approach avoids reliance on global information or complex communication networks.
+Instead, it introduces a visual anchor to address environmental boundaries, eliminating the need for explicit wall modeling.
+Through numerical simulations and experiments involving 10 independently controlled robots, the method successfully reproduces diverse collective motion behaviors.
+Figure 1: Modeling collective behaviors from optic flow and retinal cues [33]: a) View of the flock positions.
+b) Representation of flock positions after recognition at time step $t$.
+Each $i$-particle is described by its XY-coordinates and heading unit vector $\mathbf{e}_{i}$, with heading angle $\theta_{i}$.
+c) Point-of-view of particle 0, showing the relative retinal positions of each particle.
+The arrows represent the perceived optic flow, with the white arrow indicating the apparent velocity ($\mathbf{V}_{i,k}$).
+Only the azimuthal component of $\mathbf{V}_{i,k}$ is shown for retinal object $k$ as perceived by $i=0$.
+d) $\mathcal{V}(\phi)$: Binary panorama representation of the field of view of particle 0.
+e) Apparent distance $\mathcal{R}_{i}(\phi)$ for particle 0.
+f) Optic flow $\mathcal{O}_{i}(\phi)$ perceived by particle 1.
+g) Optic flow divergence $\mathcal{D}_{i}(\phi)$ perceived by particle 1.
 The article is organized as follows: Section 2.1 provides an overview of the challenges of a robotic implementation of 3A-models, with the example of the visual-based collective motion model proposed by Castro et al.
@@
 Section 2.2 describes the robotic platform and the model modifications for implementing on the 10 independently controlled robots including the introduction of the visual anchor concept.
-Section 2.4 detail the implementation of a feature matching optic flow and visual collective motion rules.
+Section 2.4 detail the implementation of a feature matching optic flow algorithm and visual collective motion rules.
 Section 2.5 describes the supplementation on original model and summarizes robotic setup, while Section 3 presents the results from simulations and robot-in-the-loop experiments.
@@
 2 Methods
-Figure 1: Model collective behaviors from optic flow and retinal cues [28].
-a) View of the flock positions.b) Representation of the flock positions after recognition at time step $t$.
-Each $i$-particle is described by its XY-coordinates and unit heading vector $\mathbf{e}_{i}$, with the heading angle $\theta_{i}$.
-c) Point-of-view of particle 0 showing the retinal relative positions of each particle.
-The arrows represent the perceived optic flow, and the white arrow indicates the apparent velocity ($\mathbf{V}_{i,k}$).
-Here, we only see the azimuthal component of the apparent velocity $\mathbf{V}_{i,k}$ by the particle i=“0” of retinal object “k”.
-d) $\mathcal{V}(\phi)$ binary panorama representation of the field of view of particle 0.
-e) Apparent distance $\mathcal{R}_{i}(\phi)$ of particle “0”.
-f) Optic flow $\mathcal{O}_{i}(\phi)$ perceived by particle 1.
-g) Optic flow divergence $\mathcal{D}_{i}(\phi)$ perceived by particle 1.
 2.1 Robotic Implementation Challenges
-Flocking models that are based on individual rules of attraction, avoidance and alignment, or better known as 3A-models, present several challenges when being implemented on a robotic platform.
-The biggest two are related to how individuals are modeled.
-The assumption of most models is that each individual is a self propelled particle, traditionally moving on a plane.
-Yet most robots are particles or restrict their movement only to a plane.
-This brings forth the first challenge, which robot kinematics should be chosen?
-It should balance following the assumptions of the model, sensing capabilities, computational capacity, size, among others.
-By exiting the numerical environment, the experimental environment are not infinite anymore, so the collective behavior should be confined which is a second challenge.
-Without this considerations flocking on robots would be as direct as a raw implementation of a established model, such as Castro et al., (2024) [28] on a the selected robotic platform (as seen on Fig.
+Flocking models based on individual rules of attraction, avoidance, and alignment—commonly referred to as 3A-models—pose several challenges when implemented on robotic platforms.
+Two significant challenges arise due to differences between theoretical models and real-world implementations.
+The first challenge is related to the abstraction of individual agents.
+Most 3A-models assume that individuals are self-propelled particles moving in a plane.
+However, many robots either deviate from the particle abstraction or lack natural constraints to planar motion.
+This raises the question: Which robot kinematics should be chosen?
+The chosen robotic platform must strike a balance between adhering to model assumptions, ensuring adequate sensing capabilities, accommodating computational constraints, and maintaining a suitable size, among other factors.
+The second challenge involves transitioning from numerical simulations to physical experiments.
+Unlike numerical environments, experimental setups are spatially constrained.
+As a result, collective behaviors must be confined within the experimental area.
+Without addressing these constraints, implementing flocking behaviors on robots would merely involve a direct adaptation of an established model, such as the one proposed by Castro et al., (2024) [33], to the chosen robotic platform (illustrated in Fig.
 1).
-The following sections will present our solution to both challenges.
+The following sections present our solutions to these challenges.
 2.2 Robotic Setup
-This section will section will present our robot setup to a robot-in-the-loop visual flocking model.
-We propose to use an underpowered, commercially available robotic platform that offloads the sensing capabilities to a base station but independently uses the information from its binary visual map to execute the model presented in the previous section.
-Figure 2: Robotic platform.
+This section describes the robotic setup for implementing a robot-in-the-loop visual flocking model.
+We propose to use educational spherical robots equipped with wireless control capabilities, while the robot’s sensing capabilities are provided by an external infrastructure.
+The external infrastructure reconstructed each binary panorama and processed the model described in Section 2.1, controlling the locomotion of each robot independently.
+Figure 2: Spherical robotic platform.
 a) Sphero Bolt®wireless educational robots.
 b) Example of the robots’ movement, c) Heading control by differential rotation on the two opposite inner wheels, d) Induced nose-up pitch during forward displacement.
-This robot-in-the-loop solution utilizes an existing educational robotic platform, the Sphero Bolt®(Fig.
+This robot-in-the-loop solution leverages an existing educational robotic platform, the Sphero Bolt®(Fig.
 2a).
-These are small, spherical robots that can be controlled in real-time via Bluetooth®and an educational app, where users can set a heading value, desired velocity, LED states, and read the onboard sensors’ values.
-Additionally, several open-source projects integrate these robots with different programming environments such as ROS, Python, or MATLAB®.
-Each robot is equipped with several sensors, including an Inertial Measurement Unit (IMU) comprising a 3-axis accelerometer, gyroscope, and magnetometer, and a light sensor.
-It also features four infrared (IR) proximity sensors on its sides, two addressable RGB LEDs (one in the front and one in the back), and an 8x8 RGB LED matrix on top.
-The Sphero’s movement is achieved via a self-balancing two-wheel enclosed system, with two non-actuated wheels maintaining constant contact with the enclosing sphere.
-This dual-motor design allows for stationary turning; however, while moving, the turning is achieved by differential motion (Fig.
-2b-c).
-The locomotion generates a pitch displacement depending on the speed and the friction coefficient between the sphere’s encapsulation and the surface (Fig.
-2d).
-This pitch is partially managed by the robot’s dual-MCU configuration: one MCU handles the communication protocol and high-level instructions, while the other manages motor control to ensure the robot achieves the desired heading and speed, along with low-level sensor readings.
-Figure 3: Experimental setup.
-a) Picture of the experimental setup, b) Example of the images taken by the camera (Daheng Imaging MER2-230-168U3C camera with 4mm c-mount lens), c) Whole robot body image segmentation, d) Individual light sources and heading recognition, e) Individual robot identification and virtualization.
-To address the lack of an onboard camera and the inability to embed one into the robots, we set up a bird’s-eye camera above the experimental area.
-This is an industrial RGB global shutter camera, the Daheng Imaging MER2-230-168U3C, with a frame rate of 168 FPS, a resolution of 1920x1200, and support for interchangeable c-mount lenses (Table 1).
-The selected lens is a 4mm, 1/8”, F2.0-12 non-distorting lens, which allows the camera to be mounted 2.3 m above the robots, covering a 2.7x2.7 m working area (approximately 38 times the length of the robots) (Fig.
+These compact, spherical robots can be controlled in real-time via Bluetooth®using an educational application, which allows users to set a heading, desired velocity, and LED states, as well as read onboard sensor data.
+Furthermore, several open-source projects extend the functionality of these robots by integrating them with programming environments such as ROS, Python, and/or MATLAB®.
+Each robot is equipped with a suite of sensors, including an Inertial Measurement Unit (IMU) composed of a 3-axis accelerometer, gyroscope, and magnetometer, as well as a light sensor.
+Additionally, the robot features four infrared (IR) proximity sensors positioned on its sides, two addressable RGB LEDs (located at the front and rear), and an 8×8 RGB LED matrix on the top.
+Locomotion is facilitated by a self-balancing two-wheel enclosed system, with two passive wheels (not shown in Fig.2) maintaining continuous contact with the spherical enclosure.
+This dual-motor configuration enables linear movement, while stationary turning is achieved via differential motion (Fig.2b-c).
+The locomotion process induces a pitch displacement dependent on the robot’s speed and the friction coefficient between the spherical enclosure and the surface (Fig.2d).
+This pitch displacement is partially mitigated by the robot’s dual-MCU architecture: one microcontroller unit (MCU) manages the communication protocol and high-level commands, while the other is dedicated to motor control.
+The latter ensures precise implementation of the desired heading and speed while handling low-level sensor readings.
+Figure 3: Experimental Setup and external infrastructure.
+a) Photograph of the experimental setup.
+b) Sample image captured by the camera (Daheng Imaging MER2-230-168U3C camera equipped with a 4 mm C-mount lens).
+c) Whole-robot body image segmentation: Segmentation of the robot’s body based on the captured image.
+d) Recognition of individual light sources and heading: Identification of the front and rear LEDs allows calculation of the robot’s heading vector.
+e) Robot identification and virtualization: Assignment of a unique identifier to each robot and visualization of their positions and orientations.
+To address the absence of onboard cameras and the inability to integrate them into the robots, an external infrastructure made of a bird’s-eye camera was positioned above the experimental area.
+The camera used is an industrial RGB global shutter model, the Daheng Imaging MER2-230-168U3C, with a frame rate of 168 frames per second (FPS), a resolution of 1920×1200 pixels, and compatibility with interchangeable C-mount lenses (Table 1).
+A 4 mm, 1/8”, F2.0–12 non-distorting lens was selected, enabling the camera to be mounted 2.3 m above the robots while covering a 2.7×2.7 m working area—approximately 38 times the length of a single robot (Fig.
 3a).
+Resolution 1920x1200
+Frame rate 168 FPS
 Shutter time 25 ms
@@
 Interface Python 3.10
-Table 1: Camera Settings.
-Each robot is assigned an individual color pattern that enables identification of the robot’s front and back.
-To maintain light connectivity between the front and back, the back half of the LED matrix is set to 50% bright white light.
-This configuration provides 16 possible combinations for identifying each robot, with the potential for expansion using additional colors.
-The bird’s-eye camera allows the isolation of each light source on top of the robots, preventing light bleeding between robots when they are in close proximity (Fig.
+Non-distorting lens 4mm (focal length), 1/8” (optical format) , F2.0-12 (aperture)
+Table 1: Daheng Imaging MER2-230-168U3C global shutter Camera Settings.
+Each robot is assigned a unique color pattern, enabling the identification of its front and back.
+To maintain visual connectivity between these sections, the back half of the LED matrix is set to emit 50% bright white light.
+This configuration supports 16 distinct combinations for robot identification, with potential for further expansion using additional colors.
+A bird’s-eye camera captures a 1920×1200 RGB image, enabling the isolation of each light source on the robots and reduces light bleeding when several robots are in close proximity (Fig.
 3b).
-This is achieved by capturing a 1920x1200 RGB image and performing whole-robot segmentation.
-This segmentation treats the three light sources as a single entity, allowing calculation of the center’s XY-coordinates, which approximate the robot’s center.
-Knowing the radius of the robots in pixels allows for cropping of $N$ images from these coordinates, each isolating a single robot.
-The cropped images are further segmented to identify the front and back LEDs, as well as the LED matrix.
-The larger area generated by the two light sources enables front and back recognition (Fig.
+Whole-robot segmentation is performed on this image, treating the three light sources on each robot as a single entity.
+This allows for the calculation of the robot’s center coordinates (X,Y), which approximate its geometric center.
+By knowing the radius of the robots in pixels, N cropped images are extracted from these coordinates, each isolating an individual robot.
+The cropped images are further segmented to recognize the front and back LEDs as well as the LED matrix.
+The larger area created by the two light sources facilitates accurate recognition of the robot’s front and back (Fig.
 3d).
-Additionally, by correlating the color from this segmentation, each robot can be uniquely identified.
-The final step is to compute the robot’s heading by tracing the vector between the center positions of the front and back LEDs.
-This process results in the identification, XY-coordinates, and heading of each robot (Fig.
+Additionally, the unique color patterns from this segmentation are used to assign a distinct identifier to each robot.
+Finally, the robot’s heading is calculated by tracing the vector formed between the center positions of the front and back LEDs.
+This computer vision processing yields the robot’s identification, X,Y-coordinates, and heading (Fig.
 3e).
-The next step, after identifying each robot and determining their location and heading, is to reconstruct the binary panorama of each robot.
+The next phase, following the identification and localization of each robot, involves reconstructing the binary panorama for each robot.
 2.3 Anchoring the collective behavior
 Figure 4: Binary Panorama Reconstruction.
-a) 2D virtual representation of the binary panorama, the red goal represents the desired location of the flock.
-b) Mixed reality point of view of the 0-particle with the omniscient red goal.
-c) Graph representation of all $\mathcal{V}(\phi,t)$.
-$\mathcal{V}(\phi,t)$ is a binary panorama representation of the field of view of particle i where the anchor is represented a different channel (red).
-This graph can be stored as a tuple list (rising and falling edges).
-In the binary panorama reconstruction, we introduce a new component to the original model [28]: a visual anchor.
+a) 2D virtual representation of the binary panorama: The red goal indicates the desired location of the flock.
+b) Mixed-reality perspective from particle 0: Displays the omniscient red goal as observed from the particle’s point of view.
+c) Graph representation of all $\mathcal{V}(\phi,t)$: $\mathcal{V}(\phi,t)$ represents the binary panorama view of the field of vision for particle i, with the anchor displayed as a distinct channel (red).
+This graph can be efficiently stored as a tuple list, representing the rising and falling edges of the binary panorama.
+In the binary panorama reconstruction, we introduce a novel component to the original model [33]: a visual anchor.
+The visual anchor acts as an attractive force directed toward a designated location, which, in this case, is the center of the arena.
+The farther an individual is from this location, the stronger the anchor’s influence, causing the individual to turn and face the anchor.
+The anchor is described by Eq.
+1, where $\mathcal{R}_{i}$ represents the apparent distance function over the binary panorama observed by particle i, and $a$ denotes the particle’s radius.
+[…]
 The inclusion of an anchor bounds the flock to a place and prevent infinite expansion.
@@
 Therefore, avoidance is implemented by providing a new piecewise constant function ($\mathcal{A}$), where attraction is inverted in situations where a particle should avoid a retinal object.
-This new function depends on the width of the retinal object, specifically the difference between its rising and falling edges ($\Delta\phi$).
-This being a indirect measurement of the apparent distance ($\mathcal{R}(\phi,t)$).
+This new function $\mathcal{A}$ depends on the width of the retinal object (optical size), specifically the difference between its rising and falling edges ($\Delta\phi$).
+Note that the apparent distance of a retinal object ($\mathcal{R}(\phi,t)$) inversely depends on this optical size $\Delta\phi$.
 We choose $\mathcal{A}(\Delta\phi)=-5$ for $\Delta\phi<\pi/10$, which is equivalent to the retinal object being closer than three particle radii ($\mathcal{R}<3a$).
@@
 The supplemented visual model equations for the collective behavior are presented in Eqs.
-1a-$f$.
+2a-$f$.
 Note that the noise $k_{\eta}$ and the time delay $\tau$ are set to 0 for robot-in-the-loop experiments, as they are parameter modeling real-world noise and uncertainty.
@@
 […]
-$\frac{\mathbf{e}_{i}(t)\times\mathbf{V}_{ik}(t)}{\mathcal{R}_{i}(t)}=-\mathcal{D}_{i}(\phi,t)\sin(\phi,t)+\overline{\mathcal{O}}_{i}(\phi,t)\cos(\phi,t).$ (1f)
-Lastly, the robotic platform introduces several critical processing factors – Bluetooth, image processing, internal Spheros control command, … – that affect the overall dynamical response of the robots and its collective behavior as a whole (see Supplementary Information for more details).
+$\frac{\mathbf{e}_{i}(t)\times\mathbf{V}_{ik}(t)}{\mathcal{R}_{i}(t)}\cdot\hat{\mathbf{z}}=-\mathcal{D}_{i}(\phi,t)\sin(\phi,t)+\overline{\mathcal{O}}_{i}(\phi,t)\cos(\phi,t).$ (2f)
+Where $\mathbf{x}_{i}(t)$ is the XY-coordinate of particle i at time t, $\mathbf{e}_{i}(t)$ is the heading vector of particle i at time t, $k_{\odot},\,k_{\parallel},\,k_{c},\,k_{\eta}$ are the gains of the attraction, parallel, anchor and noise respectively while $\omega_{\odot}(t),\,\omega_{\parallel}(t),\,\omega_{c}(t),\,\eta$ are the attraction, parallel, anchor and noise terms at timet.
+$\hat{\mathbf{z}}$ is the unitary vector that is orthogonal to the ($\hat{\mathbf{x}}$, $\hat{\mathbf{y}}$)-plane of movement of the robots.
+$\phi$ is the retinal angle in the field of view, is considered from the point of view of each particle and counter-clockwise is considered positive.
+$\mathcal{R}_{i}(\phi,t)$ is the apparent distance function at a retinal angle $\phi$ at time t for individual i.
+$b_{\epsilon}(\phi,t)$ is the blindspot function ($1+\epsilon\cos\phi$).
+$\mathbf{V}_{ik}(t)$ is the apparent velocity that an individual i perceives of a retinal object k.
+Finally $\tau$ is the delay between perception and action.
+The robotic platform introduces several critical processing factors – Bluetooth, image processing, internal Spheros control command, … – that affect the overall dynamical response of the robots and its collective behavior as a whole (see Supplementary Information for more details).
 These extra factors results in a time delay $\tau$ with a median of 600ms.
@@
 2.5 Robots-in-the-loop computer architecture
-We propose a robots-in-the-loop flocking that consists of a set of $N$ Sphero® Bolt robots, which are devoid of visual sensor and are controlled via Bluetooth®.
-Their positions and headings are calculated by processing the images from an industrial 1920x1200 RGB, 164 FPS overhead camera.
-From all their positions, a binary visual map is recreated.
-This visual map serves as the input for the collective motion model.
-This process is shown in Fig.
+We propose a robots-in-the-loop flocking framework that consists of a set of $N$ Sphero® Bolt robots, which lack visual sensors and are controlled via Bluetooth®.
+Their positions and headings are determined by processing images captured by an industrial 1920x1200 RGB, 164 FPS overhead camera.
+From these positions, a binary panorama is emulated.
+This binary panorama serves as the input to the collective motion model, as shown in Fig.
 6.
 Figure 6: General Block Diagram.
-The N flocking robots are seen by the 1920x1200px, 164fps camera.
-This image is acquired and processed to find the XY-coordinates, heading and colors for all $N$ robots.
-With these coordinates each robot can be identified.
-The XY-coordinates of the Anchor can be set at any point of the image to be integrated to the identification of the robots.
-All these coordinates are processed to reconstruct the binary panorama for each of the spherical robots.
-Each binary Panorama is the input of independent Collective motion model that process the visual panoramic information.
-All these models are ran in parallel and their output in term of heading are sent to each robot via a Bluetooth interface.
-There are two main Python programs to handle the implementation of these blocks.
-The first one (administrator.py) manages the camera communication, data extraction, image processing, object identification, and binary panorama reconstruction (Fig.
+The $N$ flocking robots are observed using a 1920x1200 px, 164 fps camera.
+The captured images are processed to determine the XY-coordinates, heading, and colors of all $N$ robots.
+Based on these coordinates, each robot is uniquely identified.
+The XY-coordinates of the Anchor can be set at any point within the image to assist in robot identification.
+These coordinates are further processed to reconstruct the binary panorama for each spherical robot.
+Each binary panorama serves as input to an independent collective motion model, which processes the visual panoramic data.
+These models operate in parallel, and their outputs, in terms of heading, are transmitted to each robot via a Bluetooth interface.
+There are two main Python programs to handle the implementation of the blocks, as shown in Fig.
+6.
+The first one (administrator.py) coordinates the camera communication, data extraction, image processing, object identification, and binary panorama reconstruction (Fig.
 3).
-The output is transferred to $N$ independent routines named PyRobot.py (the second program), which take a single binary panorama, perform the collective motion model, manage the Bluetooth connection, and send the desired heading and velocity (Fig.
+The binary panorama reconstruction is transferred to $N$ independent routines named PyRobot.py (the second program).
+Each routine processes the binary panorama of a single robot to calculate its output from the collective motion model and send the desired heading and speed via Bluetooth to that same robot (Fig.
 5).
@@
 Each PyRobot.py is a routine that can be executed either locally or remotely from the administrator.
-This setup allows for scaling the flock size to the maximum number that can fit within the camera’s recognition area.
-Note that this scaling is hardware-constrained by the Bluetooth interface capabilities of the computer.
+This setup allows a framework only bounded in the flock size to the maximum number that can fit within the camera’s recognition area and hardware-constrained Bluetooth interface capabilities of the computer.
 For Windows® machines, the theoretical maximum is seven devices per machine; in Linux, it is seven per interface.
-However, it is recommended to have a maximum of five on the same machine managing the camera and six per interface on independent machines.
+However, in practice, it is recommended to have a maximum of five on the same machine managing the camera and six per interface on independent machines.
 2.6 Metrics
@@
 These metrics are defined as follows:
-$P(t)$ $=$ $\left\|\overline{\mathbf{e}_{i}(t)}\right\|,$ (1b)
-$M(t)$ $=$ $\left\|\overline{\mathbf{y}_{i}(t)\times\mathbf{e}_{i}(t)}\right\|,$ (1c)
+$P(t)$ $=$ $\left\|\overline{\mathbf{e}_{i}(t)}\right\|,$ (2c)
+$M(t)$ $=$ $\left\|\overline{\mathbf{y}_{i}(t)\times\mathbf{e}_{i}(t)}\right\|,$ (2d)
 […]
@@
 The opacity $O(t)$ measures the “occupancy” of the visual fields: $O(t)=0$ indicates no object in the visual field, and $O(t)=1$ signifies that the entire visual field is obscured.
+To asses the performance of the phase, we quantify the number of collisions while a simulation or experiment were running.
+For an experiment, we counted how many frames presented a collision and how many robots were involved.
+For a simulation we counted collision by defining one as any time the distance between individuals was less than $2a$.
 3 Results
@@
 To make the problem dimensionless, we chose $a=1$ and $U=1$.
-We also kept the blind angle parameter and noise strength with values that lead to the most phase reproduction: $\epsilon=1$, $k_{\eta}=0.01$.
+The blind angle parameter and noise strength were set to values that enable the reproduction of most phases: $\epsilon=1$, $k_{\eta}=0.01$.
 With this approach, three dimensionless parameters remain: the strengths of anchor attraction and alignment, $k_{c}$, $k_{\odot}$, and $k_{\parallel}$.
-The anchor is a parameter aiming for a contained robotic experimentation, set practically to the minimum value that ensures the robots do not reach the arena’s border when no other law is in effect, $k_{c}=0.005$.
-The visual original model has been supplemented with three major changes impacting the phase distribution and the metric on the parameter space $(k_{\odot},k_{\parallel})$:
-  • the inclusion of the avoidance rule ($\mathcal{A}(\phi)$) (Fig. S7),
-  • the addition of an anchor law ($\omega_{c}$) (Fig. S8),
-  • the introduction of a time delay ($\tau$) (only in simulation) to reflect the time delay between the desired heading and a perceived change in heading while operating the Spheros robot, (Fig. 7).
+The anchor parameter was chosen to confine robotic experimentation, set practically to the minimum value required to prevent the robots from reaching the arena’s border when no other law is in effect, $k_{c}=0.005$.
+The original visual model has been supplemented with three significant modifications, influencing the phase distribution and metrics in the parameter space $(k_{\odot},k_{\parallel})$:
+  • inclusion of the avoidance rule ($\mathcal{A}(\phi)$) (Fig. S7),
+  • addition of an anchor law ($\omega_{c}$) (Fig.S8),
+  • introduction of a time delay ($\tau$) (simulation only) to reflect the delay between the desired heading and a perceived change in heading when operating the Spheros robot (Fig.7).
 Figure 7: Illustration of the phases observed in simulation for the supplemented visual model with the addition of avoidance, anchor, and time delay as defined in Eqs.
-1a, with $\tau=2.4[a/U]$.
+2a, with $\tau=2.4[a/U]$.
 Three phases are observed when varying the other two parameters: a) Swarming ($k_{\odot}=0.25$, $k_{\parallel}=0.04$) (see SVideo 1); b) Bistable ($k_{\odot}=0.2$, $k_{\parallel}=0.25$) (see SVideo 2); c) Milling ($k_{\odot}=0.07$, $k_{\parallel}=0.16$) (see SVideo 3).
-We now compare the original model with each consecutive addition by exploring the parameter space $(k_{\odot},k_{\parallel})\in[0,0.3]\times[0,0.3]$ on a flock of 10 individuals.
-For each parameter set, we ran 10 simulations over a long duration ($\Delta t=5000$).
-The mean values of $P$, $M$, and $O$ were determined by averaging over the last 1000 time units to ensure that transients do not influence the results.
-The outcomes of these simulations are synthesized in phase diagrams (Fig.
+We compared the original model with each successive addition by exploring the parameter space $(k_{\odot},k_{\parallel})\in[0,0.3]\times[0,0.3]$ for a flock of 10 individuals.
+For each parameter set, we ran 10 simulations over an extended duration ($\Delta t=5000$).
+The mean values of $P$, $M$, and $O$ were determined by averaging over the final 1000 time units to minimize the influence of transients.
+The simulation results are summarized in phase diagrams (Fig.
 8).
-We maintain the identification of the collective phases from the values of the polarization $P$ and milling $M$ parameters: schooling when $P>0.5$ and $M<0.5$, milling when $P<0.5$ and $M>0.5$, and swarming when $P<0.5$ and $M<0.5$.
+We classified the collective phases using the polarization $P$ and milling $M$ parameters:
+Schooling when $P>0.5$ and $M<0.5$, Milling when $P<0.5$ and $M>0.5$, Swarming when $P<0.5$ and $M<0.5$.
 Figure 8: Simulation Phase Diagrams for 10 individuals.
-The dotted line is $P=0.5$, the dashed line is $M=0.5$, and the solid line is $O=0.3$.
-A) Color code, b) Original visual model [28], c) Supplemented visual model with avoidance term, $\mathcal{A}(\mathcal{R}(\phi,t))$, d) Supplemented model with avoidance and anchor term, $\omega_{c}(t)$, e) Supplemented visual model with avoidance, anchoring and perceived time delay ($\tau$=2.4[a/U]).
+The dotted line indicates $P=0.5$, the dashed line indicates $M=0.5$, and the solid line indicates $O=0.3$.
+a) Color code, b) Original visual model [33], c) Supplemented visual model with avoidance term, $\mathcal{A}(\mathcal{R}(\phi,t))$, d) Supplemented model with avoidance and anchor term, $\omega_{c}(t)$, e) Supplemented visual model with avoidance, anchoring, and perceived time delay ($\tau=2.4[a/U]$).
 3.2 Phase Diagrams of the Supplemented Model in Simulation
-We now compare the model with each consecutive addition and the robotic implementation, exploring the parameter space $(k_{\odot},k_{\parallel})\in[0,0.3]\times[0,0.3]$ for a flock of 10 independently controlled robots.
-For the robotic implementation, the dimensionless parameters are set to match the radius in pixels and velocity in radius per second, resulting in $a=1$, with the other parameters remaining the same ($\epsilon=1$, $k_{\eta}=0$, and $k_{c}=0.005$).
-For each parameter set, we ran 5 robotic experiments over a long duration (240sec or $\Delta t\approx 8500$).
-The mean values of $P$, $M$, and $O$ were determined by averaging over the last 1000 time units to ensure that transients do not influence the results.
-The outcomes of these experiments are synthesized in phase diagrams (Fig.
-9a).
-An example of swarming is shown in Fig.
-9b, of partial schooling for a bistable phase in Fig.
-9c, and of milling in Fig.
-9d.
+We further compared the supplemented model and its robotic implementation by exploring the parameter space $(k_{\odot},k_{\parallel})\in[0,0.3]\times[0,0.3]$ for a flock of 10 independently controlled robots.
+For robotic experiments, dimensionless parameters were adjusted to match the radius in pixels and velocity in radius per second, resulting in $a=1$, with other parameters remaining constant ($\epsilon=1$, $k_{\eta}=0$, and $k_{c}=0.005$).
+For each parameter set, five robotic experiments were conducted over an extended period (240 sec or $\Delta t\approx 8500$).
+The mean values of $P$, $M$, and $O$ were calculated by averaging over the final 1000 time units to eliminate transient effects.
+The experimental results are summarized in phase diagrams (Fig.9a).
+Examples of swarming, partial schooling in a bistable state, and milling are presented in Fig.9b, Fig.9c, and Fig.9d, respectively.
+Note that the bistable state has been observed in fish experiments ([46]).
 The phase diagram of the original model [33] showcases the reproduction of the main three phases (swarming, schooling, and milling), with the area above $k_{\odot}>0.15$ and $k_{\parallel}>0.15$ being bistable (Fig.
 8a).
-A bistable phase oscillates between milling and schooling (Fig.
+A bistable state is when the flock oscillates between milling and schooling (Fig.
 8b).
@@
 3.3 Comparison with the phase diagram of the Spheros’ robots-in-the-loop
-For the robotic implementation (see Fig.
-9), we found that the milling phase remains fairly stable, appearing for high alignment ($k_{\parallel}>0.15$) across most attraction values ($k_{\odot}<0.16$).
+During the robotic implementation (see Fig.
+9), the milling phase remains fairly stable, appearing for high alignment ($k_{\parallel}>0.15$) across most attraction values ($k_{\odot}<0.16$).
 The swarming phase is constant for low alignment ($k_{\parallel}<0.06$) or high attraction ($k_{\odot}>0.17$).
@@
 4 Discussion
-The collective motion model with the robotic platform reproduces several collective patterns such as swarming, milling, and a bistable phase.
-For the first time to the best of our knowledge, a visual flocking on spherical robots shows a phase diagram that matches almost perfectly that of the simulation with the same visual model (Fig.
-8e vs Fig.
-9a).
-The mean squared error (MSE) between the phase diagram images of robots-in-the-loop and supplemented visual model results is 2.48% (for more details see Fig.
+In this article, we presented a robot-in-the-loop approach to a visual collective motion model.
+The robots used are practical to deploy and commercially available: the Sphero BOLT (Fig.
+2).
+The Sphero BOLTs are sphere-like robots with a transparent shell that encloses all electronics, sensors, motors, batteries, and LEDs in a sealed design.
+However, due to this sealed design, integrating vision required an external solution.
+We opted for an external infrastructure made of overhead camera system to recognize the robots, enabling the reconstruction of each robot’s visual field.
+Each robot was assigned a unique light pattern, visible through the transparent shell, to facilitate position and heading recognition via the overhead camera.
+(Fig.
+3).
+The noise effect
+The collective motion model presented by Castro et al.
+(2024) [33] and the supplemented version introduced in this article, incorporated noise as a component during the computation of angular velocity model output which is directly used to actuate the robot.
+Introducing the noise on the actuation signal accounted for the noise that errors on the heading recognition might introduce, but did not accounted for recognition errors of the position.
+And, by extension, the noise that those errors might introduce on the robots’ panorama reconstruction hence on the measurement of the “early vision” cues (apparent distance, optical position, and optic flow signals on Fig.
+1).
+The effect of noise on these “early vision” cues can be particularly significant for the alignment term in the model.
+Alignment is calculated using optic flow and apparent distance (Eq.
+2d), making it three times more sensitive to measurement noise than attraction that only uses apparent distance (Eq.
+2c).
+Fig.
+S1 illustrates the combined effect of these recognition errors in open loop, showing that the mean error is approximately 6%.
+Consequently, the attraction term inherently has a 6% error, while alignment error reaches around 15%.
+Despite this, the simulation and experimental robotic results align on a macro level as shown in Fig.
+9.
+For comparison, the mean squared error (MSE) between the phase diagram images from robots-in-the-loop and supplemented visual model results is as low as 2.48% (for more details, see Fig.
 S2).
-Such implementation therefore bridges the gap between collective model simulation and real-world experiments, allowing for expansion to $N$ individuals with low complexity.
-A model based on visual information enables a mixed-reality approach where there can be interaction between the flock and the environment without adding complexity to the model or sensors available to each robot.
-Future works with these spherical robots could include an explicit channel for avoidance also using optic flow, a moving anchor, a replacement of the anchor by another particle without adding a new channel, and an expansion of the flock size to 20 or 40 individuals for a closer-to-nature implementation of these visual collective models.
+Effect of collisions
+We analyzed the number of collisions throughout the evolution of the flock during the examples of swarming, bistability, and milling.
+If a collision occurred during the robotic experiment or simulation, we counted how many individuals participated on the collision (as seen on Fig.
+S9).
+In simulation, a collision is detected when the distance between individuals is less than or equal to $2a$.
+In the robotic experiment, a collision is detected when the recognized object becomes much larger than a single Sphero.
+These collisions are calculated throughout the simulation and the robotic experiment, and are presented as additional information on the state of the flock (Fig.
+S10).
+We found that, on average, in one video for each phase, the flock in simulation spent 14% of their time experiencing collisions.
+The robots exhibited a collision rate of 33% of the time (on average, three times more than the simulation).
+Interestingly, most simulation collisions were between two particles, and instances of collisions involving more than two particles were rare (only seen on Fig S9-b).
+In contrast, with the robots, most collisions caused a chain reaction, leading other particles to collide with the stopped particles (as seen on Fig S9-d).
+The most collisions occurred during the bistable state, where the simulation had a 21% collision rate compared to 36% for the robots.
+In both cases, the fewest collisions occurred during the milling phase.
+The simulation only showed a 6% collision rate, while the robots exhibited a 30% collision rate.
+The stability of the milling phase is also seen on the minimum inter-distance that, on simulation and on the robots, tends to stabilize around t=1200.
+The timeline of this phase can be seen on Fig S9-c for simulation and Fig S9-f for robots.
+The anchor’s effect
+One notable modification to the model was the introduction of an attractive force towards the center of the area of interest, referred to as the anchor (Eq.
+1).
+As shown in Fig.
+8c, the main phase is schooling.
+However, when the anchor is introduced (Fig.
+8d), schooling disappears.
+We hypothesize that the anchor, acting as a force that causes individuals to turn towards the center when they are far from it, challenges the formation of schooling.
+Notably, schooling is typically observed during flock translation, such as animal migration, where the flock’s center of mass moves along the flock’s polarized direction.
+If the model bounds a flock with a wall, a schooling flock would inevitably interact with the wall.
+In contrast, our model introduces a force directing individuals toward the a desired location, the center of the flock.
+This means a schooling flock, while temporarily moving away, would eventually return toward the selected point.
+Moreover, the anchor acts as an unbiased attraction force, causing individuals to turn either clockwise or counterclockwise.
+During this turning, the flock depolarizes, leading to phase changes.
+As a result, any observed schooling is temporary.
+The delay’s effect
+Another significant factor contributing to the expansion of milling is the delay.
+A delay between recognition and actuation has been documented to influence collective motion.
+[47, 48, 49, 50].
+If the anchor interacts with the flock with an attractive force, the delays does it with an alignment force.
+Because the delay, the particles tend to align with a past velocity.
+In other words, the polarization perceived by the particles is delayed, which leads to longer convergence on the 2 coordinated phases, only schooling or milling.
+This longer convergence time that the coordinated phase takes to stabilize creates opportunities for other phases to replace the current coordinated phase.
+Particularly the swarming phase expands because the flock itself maintains a uncoordinated cohesion on high attraction or low alignment (Fig.
+8d).
+For the case of the bistability region, it expands on the borders of the coordinated phases, as the delay causes the flock to oscillate between milling and schooling phases more often.
+Comparison between phase diagrams
+The comparison between the robotic (Fig.
+9a) and the simulated phase diagrams (Fig.
+8d) only differs for the small area of polarized bistability.
+The low-attraction region differs because collisions are more relevant in the experimental case than in simulation, where avoidance was included only in the attraction mechanism (low attraction equals low avoidance).
+In the high-attraction region, the small difference is attributable to the significant noise present in the alignment rule, caused by the compounding effect of optic flow calculations based on retinal location and apparent distance.
+All three primary effects of the robotic implementation expands the area of a stable milling phase.
+5 Conclusion
+Regardless of these implementation challenges, the collective motion model applied on the robotic platform reproduces several collective patterns, such as swarming, milling, and a bistable phase.
+To the best of our knowledge, for the first time, a visual flocking implementation using spherical robots produces a phase diagram that matches almost perfectly with that of the simulation counterpart using the same visual model.
+Such an implementation bridges the gap between collective model simulation and real-world experiments, allowing for a framework independent of the $N$ individuals belonging to the flock.
+Future Work
+For future work, it would be interesting to implement an obstacle avoidance rule based on optic flow.
+With the divergence of the optic flow, potential collisions could be better avoided.
+It would also be valuable to explore how the presence of the most influential retinal object could affect the phase distribution or evaluate the current model to identify the average location of the most influential retinal object.
+Another interesting approach would be to introduce a true blind spot where there is no sensing.
+Potential applications include the evaluation of a moving anchor or transforming the concept of the anchor into a “guarding dog” individual—one individual that moves around and keeps the flock in the desired position—or to include the anchor as a static individual.
+Another interesting path would be the implementation to 20 or 40 individuals, or even incorporating a mixed-reality option with virtual individuals that interact with the robots to create flocks of hundreds.
 Acknowledgements
+We would also like to thank the two reviewers for their valuable suggestions and comments.
 The participation of D.C. in this research was made possible by joint PhD grant from Aix-Marseille University.
@@
 Data Availability
-Data acquired during the experiments and simulations are available online [41, 42]
+Data acquired during the experiments and simulations are available online [51]
 Video Accessibility